Solve the following system of equations. 2x+4y = 8 3x + 4y = 16 x = 0 Olo X y = -1
Answer #1

2ҳ+ Иy = & 2х + чу = 16 — Ф — subtract - eqn from eq Зҳ + Чу = 16 Эх + ну 28 х = 8. Put value of n in egn - 2 (8) + чу = 8 -